OpenStudy (anonymous):

what is 56/78 times 65/88 in simplest form?

zepdrix (zepdrix):

Hey there! \(\Large\bf \color{#CC0033}{\text{Welcome to OpenStudy! :)}}\)

zepdrix (zepdrix):

\[\Large\rm \frac{56}{78}\cdot\frac{65}{88}\]

zepdrix (zepdrix):

Hmm I think it will be easier if we try to do some factoring BEFORE multiplying.

zepdrix (zepdrix):

88 can be written as 8 times 11, yes? \[\Large\rm \frac{56}{78}\cdot\frac{65}{8\cdot11}\]